### v1

@@ 3 Data and Stylized Facts
This section describes the data and preliminary empirical analysis. Section 3.1 outlines the main data sources. Section 3.2 discusses how we adapt the baseline model to the data to construct key variables. Section 3.3 details the sample selection and provides summary statistics. Finally, Section 3.4 presents evidence testing the model’s predictions.
@@ 3.1 Data Sources
Our main dataset is the U.S. Census Bureau’s Linked/Longitudinal Firm Trade Transaction Database (LFTTD), which covers the universe of U.S. import transactions from 2001 to 2018. Each observation corresponds to a shipment from a foreign exporter to a U.S. importer and includes the transaction date, product classification at the 10-digit Harmonized System (HS10) level, FOB import value in U.S. dollars, physical quantity, transportation mode, and country of origin. Exporters are identified using a manufacturer ID (MID) constructed by the Census Bureau from the exporter’s name, street address, city, and country. [^17] The MID combines the country code, (elements from) the firm name, city, and address (Kamal and Monarch, 2018). Because the algorithm is not standardized, it may generate inconsistent identifiers due to misspellings or minor location changes, leading to one firm having multiple MIDs or several firms sharing one. Following Kamal and Monarch (2018), we construct a robustness version that truncates location fields to improve consistency. Our baseline uses the full MID, and results are robust to this alternative.
To focus on arm’s-length trade, we exclude related-party transactions from the baseline sample. The LFTTD includes a related-party indicator based on a mandatory field in U.S. Customs forms, flagging relationships with ownership stakes of at least five percent. While widely used, this measure may misclassify firms due to its reliance on self-reporting and a low reporting threshold (Ruhl, 2015). To improve accuracy, we construct an alternative indicator using ORBIS, which provides firm-level cross-border ownership links. We merge ORBIS to the LFTTD as described in Appendix C.1.
We supplement the transaction-level data with information on statutory U.S. import tariffs introduced during the 2018 trade war. We use the dataset from Fajgelbaum et al. (2020), which records the timing, product coverage, and country-specific scope of these measures at the HS8–month level. The tariffs averaged 25 percentage points and were imposed on top of existing rates, targeting selected goods. They were implemented in phases over the course of the year, beginning with imports from China and later expanding to goods from other trade partners, including Canada, Mexico, and the European Union. Tariff changes are annualized based on the number of months each measure was in effect.
@@ 3.2 Measuring Key Variables of the Model
and $\alpha_{jh}\in(0,1)$ denotes the (observed) Cobb-Douglas share of input $h$ in firm $j$’s total imports of foreign intermediates. This formulation implies that the elasticity of the importer’s marginal cost with respect to the price of foreign input $h$ is $\frac{d\ln c_{j}}{d\ln p_{jh}^{f}}=\alpha_{jh}\gamma\in(0,1]$.
@@
In contrast, the importer’s buyer share is constructed as $x_{ijh}\equiv\frac{q_{ijh}}{\sum_{k\in\mathcal{Z}_{i}^{h}}q_{ihk}}$, where $\mathcal{Z}_{i}^{h}$ is the set of all U.S. importers buying product $h$ from exporter $i$. Since our dataset only includes U.S. importers, we assume that exporter $i$ operates product- and destination-specific production lines. Under this assumption, the denominator of $x_{ijh}$, which captures the total quantity of product $h$ from exporter $i$, includes only those sold to U.S. buyers. This restriction reflects a data limitation, as we do not observe importer destinations beyond the U.S. and thus cannot account for the full set of an exporter’s buyers. [^18] To address the possibility that the importer’s buyer share $x_{ij}$ may be overstated due to unobserved sales to other destinations, we replicate the analysis using only exporters from Canada and Mexico. These countries direct the majority of their exports to the U.S. (71% for Canada and 73% for Mexico in 2019), making the assumption of destination-specific production less restrictive. Reassuringly, our estimates remain stable in this subsample. Full results are available upon request.
@@ 3.3 Sample Construction and Summary Statistics
We apply a series of restrictions to the LFTTD to align the empirical sample with the model’s focus on decentralized bargaining over intermediate inputs. Full details are provided in Appendix C.2. The selection criteria are designed to ensure that we observe relationship-level price changes, exclude related-party transactions, and maintain sufficient variation in buyer–supplier matches for identification purposes. Appendix Table C.1 summarizes how the sample evolves with each restriction across key dimensions.
We begin by restricting the sample to importer–exporter–product triples observed in two consecutive years, allowing us to compute relationship-level price changes, which are central to our analysis. While this requirement eliminates over half of the raw matches, it retains a substantial share of trade, covering roughly 88% of import value between 2001–2016 and 80% between 2017–2018, highlighting the importance of repeated relationships.
Next, we restrict the sample to capital and intermediate inputs by excluding products classified as consumption goods under the Broad Economic Categories (BEC) system. This step reduces the number of suppliers and relationships, as shown in Table C.1.191919As a robustness check, we also consider a broader sample that includes consumption goods. Appendix Table C.2 reports the composition of this sample and Table C.3 reports the corresponding summary statistics. These statistics are broadly similar to the baseline, indicating that including consumption goods does not significantly alter the data composition. To further mitigate measurement error, we apply three filters: (i) exclude transactions involving energy goods, (ii) drop observations with unit values outside the 1st–99th percentile within product, and (iii) remove transactions with absolute log price changes above four.
We then exclude related-party transactions, which are less likely to reflect decentralized bargaining and may involve internal pricing strategies such as transfer pricing. [^20] Bernard et al. (2006) document that prices in related-party trade differ systematically from arm’s-length transactions, with lower average prices and distinct pass-through behavior. In our baseline definition, a buyer–supplier pair is considered related if ORBIS identifies a shared corporate parent. [^21] We retain all observations not flagged as related in either ORBIS or the LFTTD. This approach preserves sample size while ensuring that limited ORBIS coverage does not unduly constrain the selection. See Appendix C.1 for details. For robustness, we consider two alternative definitions: one based solely on the LFTTD’s related-party flag, and another that combines this flag with ORBIS data identifying the U.S. importer as a multinational (domestic or foreign-owned).
Lastly, we impose restrictions based on our identification strategy. Since the latter relies on observing the variation in prices across U.S. buyers for the same supplier-product-year combination, we restrict our sample to supplier–product pairs in which the supplier transacts with at least two U.S. buyers in consecutive years. Table C.1 reports that after these restrictions, the sample accounts for approximately \$160 billion in import value and 250 thousand buyer-supplier-product-year combinations between 2017-2018.
@@
Table 1 reports summary statistics for our final sample. Panel A shows that the concentration of importers and exporters is substantial. On average, an exporter supplies 32% of an importer’s total imports of a given HS10 product, with a median share of 15%. The average buyer share is lower, at 25%, with a median of 10%. The two shares are highly dispersed and largely uncorrelated, with a correlation of 0.04.
@@
Variable Mean Std. Dev. P25 Median P75
Panel A: Characteristics of Trade Relationships
$s_{ijh}$: Supplier share 0.32 0.35 0.03 0.15 0.57
$x_{ijh}$: Buyer share 0.25 0.29 0.02 0.10 0.40
 Relationship length (product $h$) 4.00 2.80 2.50 3.50 5.50
 Relationship length (all products) 4.80 3.30 2.50 4.50 6.50
 # Transactions (product $h$) 120 1100 6.50 16 50
 # Transactions (all products) 360 3000 11 36 140
 # Products per pair 3.80 7.30 1.50 2.50 4.50
 Multi-HS10 dummy 0.59 0.49 0.00 1.00 1.00
 # Suppliers per buyer (HS10) 1.80 3.20 1.50 2.50 5.50
 Buyer tenure (all products) 9.90 5.00 6.50 10.00 14.00
 Buyer tenure (product $h$) 6.90 4.40 3.50 6.50 10.00
 # Buyers per supplier (HS10) 3.20 3.90 2.50 3.50 7.50
 Supplier tenure (all products) 8.00 4.60 4.50 8.50 12.00
 Supplier tenure (product $h$) 6.40 4.00 3.50 6.50 9.50
 Corr. between $s_{ijh}$ and $x_{ijh}$ 0.041
Panel B: Prices
$\log p$ (pre-duty) 3.50 2.80 1.40 3.10 5.40
$\log p$ (pre-duty, excl. charges) 3.40 2.80 1.30 3.00 5.40
$\log p^{\text{duty}}$ (post-duty) 3.50 2.80 1.40 3.10 5.40
Notes: This table reports summary statistics for the estimation sample used in the empirical analysis. The data span 2001–2016 and include importer–supplier–product matches observed in two consecutive years. The sample excludes consumption goods (based on BEC), energy products, statistical outliers, and related-party trade, and is restricted to suppliers trading with at least two U.S. buyers in two consecutive years. This sample corresponds to the “+ Supplier Multi-Buyer” row in Panel B of Table C.1. Columns report the mean, standard deviation, and selected quantiles (25th, 50th, 75th percentile) for each variable. Prices in Panel B are log unit values (FOB value over quantity), with variants including charges or duties. $s_{ijh}$ denotes exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ denotes buyer $j$’s share in exporter $i$’s U.S. exports of the same product. Relationship length and tenure are in years; concentration is measured at the HS10–year level. Counts of buyers, suppliers, and origin countries are per product per firm. Statistics are based on confidential LFTTD data and rounded to four significant digits per U.S. Census Bureau Disclosure Guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Long-term relationships are a hallmark of intermediate input trade (Antràs and Chor, 2013; Monarch, 2022). In our data, importer–exporter pairs trade the same product for an average of four years and remain connected across all products for nearly five years. Pairs transact frequently (median: 16 times per product), often spanning multiple products (mean: 3.8), and typically involve durable links, with average tenures ranging from 7 to 10 years.
Panel B of Table 1 reports descriptive statistics for three bilateral (log) price measures based on FOB unit values: baseline pre-duty prices, pre-duty prices excluding insurance and other ancillary charges, and post-duty prices. All three exhibit wide dispersion across importer–exporter–product matches, with interquartile ranges exceeding four log points.
@@
To assess the sources of this variation, we perform a variance decomposition exercise, as described in Appendix D.1. Table D.1 shows that product-year fixed effects explain about 50% of the total variance, while match-specific residuals account for 4%. Crucially, when we focus only on variation within supplier–product–year cells, 77% of the remaining dispersion is explained by match-specific factors. This pattern holds across price definitions, underscoring the importance of relationship-specific forces in pricing. [^22] See Fontaine et al. (2020) for related evidence in French data.
3.4 Test of Model Predictions
We now examine how the comovements between markups, pass-through, and bilateral market shares align with the model’s predictions in Propositions 2 and 4. While not a formal test, this analysis provides supporting evidence for the mechanisms emphasized in the theory. Section 4 then develops a structural approach to quantify these forces more directly.
3.4.1 Test of Proposition 2: Markups and Bilateral Market Shares
We begin by testing Proposition 2, which predicts that with two-sided market power ($\phi\in(0,1)$ and $\theta<1$), bilateral markups increase with the exporter’s supplier share ($s_{ijht}$) and decrease with the importer’s buyer share ($x_{ijht}$). Since markups are not observed, we use log prices, equal to log markups plus log marginal costs, and include supplier–product–time fixed effects to absorb cost variation and isolate the markup component.
We estimate the following specification:
@@ 3.4.1 Test of Proposition 2: Markups and Bilateral Market Shares
where the coefficients of interest are $\alpha_{s}$ and $\alpha_{x}$, which we expect to be positive and negative, respectively.
To address endogeneity concerns, we construct leave-one-out instruments that isolate variation in market structure plausibly exogenous to the pricing decision of a given buyer–supplier pair. Specifically, we use the average supplier share (excluding $i$) among other buyers of exporter $i$ (excluding $j$) to instrument for $s_{ijh}$. Similarly, to instrument for $x_{ijh}$, we use the average buyer share (excluding $j$) among other suppliers to importer $j$ (excluding $i$).
Table 2 shows the results. Columns (1)–(2) include exporter ($\text{FE}_{i}$), importer ($\text{FE}_{j}$), and product–year ($\text{FE}_{ht}$) fixed effects. Column (3)–(4) replaces exporter fixed effects with exporter–product–year ($\text{FE}_{iht}$) to account for unobserved marginal costs. Columns (5) and (6) further replace importer fixed effects with importer–product–year ($\text{FE}_{jht}$) to capture buyer-specific demand shocks. All regressions control for the relationship duration, measured as years since the first shipment of product $h$ between $i$ and $j$.
@@
Notes: This table reports OLS and IV estimates of equation (3.1), where the dependent variable is the log FOB unit value of product $h$ imported by buyer $j$ from supplier $i$ in year $t$. Columns alternate between OLS and IV specifications. All regressions control for log relationship length (in years) within HS10 products. Columns (1)–(2) include buyer ($\text{FE}_{j}$), supplier ($\text{FE}_{i}$), and product-year ($\text{FE}_{ht}$) fixed effects. Columns (3)–(4) use supplier–product–year ($\text{FE}_{iht}$) and buyer ($\text{FE}_{j}$) fixed effects. Columns (5)–(6) include fully interacted buyer–product–year and supplier–product–year fixed effects ($\text{FE}_{jht}$, $\text{FE}_{iht}$), flexibly controlling for sourcing and pricing patterns. IV estimates (even-numbered columns) use leave-one-out instruments: $s_{ijht}$ is instrumented with the average share of other suppliers across buyers of $i$ (excluding $j$), and $x_{ijht}$ with the average share of other buyers across suppliers to $j$ (excluding $i$). Because the model includes multiple endogenous regressors, we report both first-stage and conditional F-statistics from Sanderson and Windmeijer (2016), which assess instrument strength for each endogenous regressor conditional on the others, addressing limitations of standard first-stage tests in multi-equation IV settings. Standard errors are robust. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
We estimate $\alpha_{s}>0$ and $\alpha_{x}<0$, with both coefficients statistically and economically significant. A positive $\alpha_{s}$ indicates oligopoly power, requiring exporter bargaining power ($\phi<1$), while a negative $\alpha_{x}$ reflects oligopsony power, requiring importer bargaining power ($\phi>0$) and upward-sloping supply ($\theta<1$). These findings are consistent with the model’s core assumption of two-sided market power, i.e., $\phi\in(0,1)$ and $\theta<1$.
3.4.2 Test of Proposition 4: Pass-Through and Bilateral Market Shares
We next test Proposition 4, which links tariff pass-through to the importer’s buyer share. Under decreasing returns ($\theta<1$), the model predicts that pass-through declines with the buyer share via the cost channel, providing a direct test for $\theta<1$.
For this analysis, we focus on 2017–2018, when U.S. imports experienced sharp and unanticipated tariff increases under the Trump administration. We estimate the following regression specification:
$\Delta\ln p_{ijht}=$ $\alpha_{0}+\alpha_{1}\,\Delta\ln(1+\tau_{cht})+\alpha_{s}\,\Delta\ln(1+\tau_{cht})\cdot s_{ijh,t-1}+\alpha_{x}\,\Delta\ln(1+\tau_{cht})\cdot x_{ijh,t-1}$
$+\alpha_{2}\,s_{ijh,t-1}+\alpha_{3}\,x_{ijh,t-1}+\textbf{X}_{ijht}\boldsymbol{\gamma}\,+\mathbf{FE}+\epsilon_{ijht}.$ (3.2)
where we use the change in the duty-exclusive price as the dependent variable, defined as $\Delta\ln p_{ijht}=\Delta\ln p_{ijht}^{\text{duty}}-\Delta\ln(1+\tau^{\text{app}}_{cht})$, where $\tau^{\text{app}}_{cht}$ is the applied ad-valorem tariff. This transformation isolates price changes net of applied duties, mitigating measurement error.
@@ 3.4.2 Test of Proposition 4: Pass-Through and Bilateral Market Shares
We estimate two specifications. The first includes product–time ($\text{FE}_{ht}$) and exporting country–sector ($\text{FE}_{cs}$) fixed effects, following standard practice in the pass-through literature. This is our baseline. The second is more demanding, adding importer–time ($\text{FE}_{jt}$) and exporting country–time ($\text{FE}_{ct}$) fixed effects.
@@
$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.026 0.042
@@
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.050 0.054 0.053
@@
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.403 -0.403 -0.411
@@
Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$. Columns (2) and (6) interact tariffs with the log of relationship longevity, measured as the number of years that buyer $j$ and supplier $i$ have transacted in product $h$. Columns (3) and (5) interact tariffs with the lagged supplier share, $s_{ijht-1}$, defined as supplier $i$’s share in buyer $j$’s imports of product $h$. Columns (4) and (5) interact tariffs with the lagged buyer share, $x_{ijht-1}$, defined as buyer $j$’s share in supplier $i$’s exports of product $h$. All regressions include product–year and exporter country–sector fixed effects ($\text{FE}_{ht}+\text{FE}_{cs}$). Controls include: (i) $\ln\text{longevity}_{ijht}$; (ii) $\Delta\ln q_{i(-j)ht}$, the change in exporter $i$’s total sales of $h$ to U.S. buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the weighted average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights. Standard errors are clustered at the HS8 product and exporter-country level. The sample corresponds to the ”+ Supplier Multi-Buyer” definition in Table C.1. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. See Table D.11 for results using an alternative definition of arm’s-length trade based on LFTTD. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).

### v2

@@ 3 Data and Stylized Facts
This section describes the data and preliminary empirical analysis. Section 3.1 outlines the main data sources. Section 3.2 discusses how we adapt the baseline model to the data to construct key variables. Section 3.3 details the sample selection and provides summary statistics. Finally, Section 3.4 presents reduced-form evidence on prices and pass-through through the lens of the model.
@@ 3.1 Data Sources
Our main dataset is the U.S. Census Bureau’s Linked/Longitudinal Firm Trade Transaction Database (LFTTD), which covers the universe of U.S. import transactions from 2001 to 2018. Each observation corresponds to a shipment from a foreign exporter to a U.S. importer and includes the transaction date, product classification at the 10-digit Harmonized System (HS10) level, Free on Board (FOB) import value in U.S. dollars, physical quantity, transportation mode, and country of origin. Exporters are identified using a manufacturer ID (MID) constructed by the Census Bureau from the exporter’s name, street address, city, and country. [^19] The MID combines country code, elements of the firm name, city, and address (Kamal2018). Because the algorithm is not standardized, identifiers may be inconsistent: one firm may have multiple MIDs or multiple firms may share one. Following Kamal2018, we construct a robustness measure that truncates location fields to improve consistency. Results are robust.
To focus on arm’s-length trade, we exclude related-party transactions from the baseline sample. The LFTTD includes a related-party indicator based on a mandatory field in U.S. Customs forms, flagging relationships with ownership stakes of at least five percent. While widely used, this measure may misclassify firms due to its reliance on self-reporting and a low reporting threshold (kimintrafirmAERPP). To improve accuracy, we construct an alternative indicator using ORBIS, which provides firm-level cross-border ownership links. We merge ORBIS to the LFTTD as described in Appendix C.1.
We supplement the transaction-level data with information on statutory U.S. import tariffs introduced during the 2018 trade war. We use the dataset from fajgelbaum2020return, which records the timing, product coverage, and country-specific scope of these measures at the HS8-month level. The tariffs averaged 25 percentage points and were imposed on top of existing rates, targeting selected goods. They were implemented in phases over the course of the year, beginning with imports from China and later expanding to goods from other trade partners, including Canada, Mexico, and the European Union. Tariff changes are annualized based on the number of months each measure was in effect.
@@ 3.2 Measuring Key Variables of the Model
and $\alpha_{jh}\in(0,1)$ denotes the (observed) Cobb-Douglas share of input $h$ in firm $j$’s total imports of foreign intermediates. This formulation implies that the elasticity of the importer’s marginal cost with respect to the price of foreign input $h$ is $\frac{d\ln c_{j}}{d\ln p_{jh}^{f}}=\alpha_{jh}\gamma\in(0,1]$. We also note that allowing this elasticity to vary across importer-product pairs in turn makes $\eta$ in equation (2.13) to vary across the same dimension, $\eta_{jh}$.
@@
In contrast, the importer’s buyer share is constructed as $x_{ijh}\equiv\frac{q_{ijh}}{\sum_{k\in\mathcal{Z}_{i}^{h}}q_{ihk}}$, where $\mathcal{Z}_{i}^{h}$ is the set of all U.S. importers buying product $h$ from exporter $i$. Since our dataset only includes U.S. importers, we assume that exporter $i$ operates product- and destination-specific production lines. Under this assumption, the denominator of $x_{ijh}$, which captures the total quantity of product $h$ from exporter $i$, includes only those sold to U.S. buyers. This restriction reflects a data limitation, as we do not observe importer destinations beyond the U.S. and thus cannot account for the full set of an exporter’s buyers. [^20] To address potential overstatement of buyer share $x_{ij}$ due to unobserved exports to other destinations, we replicate the analysis for Canada and Mexico, which send most exports to the U.S. (71% and 73% in 2019). Estimates are similar in this subsample. Full results available upon request.
@@ 3.3 Sample Construction and Summary Statistics
We apply a sequence of restrictions to the LFTTD to construct the estimation sample. Some restrictions are required for identification and counterfactuals; others align the baseline sample with the model’s environment. Appendix C.2 provides details, and Appendix Table C.1 traces the sample under each restriction.
Essential restrictions.
Three restrictions are essential. First, we require importer-exporter-product triplets to be observed in two consecutive years, a necessary restriction for bilateral price to be computed. Second, because identification of the bargaining and returns-to-scale parameters relies on cross-sectional price dispersion across importers sourcing the same exporter-product-year, we restrict the sample to exporter-product pairs that transact with at least two U.S. buyers in consecutive years. Third, to limit measurement error and ensure that unit values more closely reflect prices rather than compositional changes within HS10 products, we follow standard practice and (i) exclude energy goods, (ii) drop unit values outside the 1st-99th percentile within product, and (iii) remove observations with absolute log price changes exceeding four.
Other restrictions.
We impose two additional restrictions to align the data with the model. First, we exclude related-party transactions, as prices in such relationships may reflect forces, such as transfer pricing, outside our framework (bernard2006transfer). In the baseline, pairs are classified as related if ORBIS identifies a common corporate parent; alternative definitions based on the LFTTD related-party flag or multinational status yield similar results.
Second, we focus on capital and intermediate inputs by excluding consumption goods under the Broad Economic Categories (BEC) classification. This reflects the model’s focus on supply chains, where lock-in from relationship-specific investments limits short-run scalability and is more pronounced upstream (e.g., Nunn2007). In robustness exercises, however, we show that the main results extend to the full set of imported goods.
After all restrictions, the baseline sample covers approximately \$160 billion in import value and 250,000 buyer-supplier-product-year observations in 2017-2018. This corresponds to about 20% of total U.S. imports by value and roughly 15% of buyer-supplier-product triplets. Although not representative of aggregate U.S. trade, the sample isolates repeated, arm’s-length production relationships where relationship-level price changes can be credibly measured and the model’s mechanisms credibly identified.
@@
Table 1 reports summary statistics for our final sample. [^21] Appendix Table C.2 provides additional statistics. Appendix Table C.3 and Table C.4 report sample composition and summary statistics for the pooled sample including all BEC categories. Panel A documents substantial concentration on both sides of the market. On average, an exporter supplies 32% of an importer’s total imports of a given HS10 product, with a median share of 15%. The average buyer share is lower, at 25%, with a median of 10%, likely reflecting the exclusion of cases with buyer share equal to one. The two shares are highly dispersed and essentially uncorrelated (correlation 0.04).
@@
Variable Mean Median P25–P75
Panel A: Trade Relationships
$s_{ijh}$: Supplier share 0.32 0.15 0.03–0.57
$x_{ijh}$: Buyer share 0.25 0.10 0.02–0.40
Relationship length 4.0 3.5 2.5–5.5
Relationship length (all products) 4.8 4.5 2.5–6.5
# Transactions 120 16 6.5–50
# Products per pair 3.8 2.5 1.5–4.5
Buyer tenure 6.9 6.5 3.5–10.0
Supplier tenure 6.4 6.5 3.5–9.5
Panel B: Prices (log unit values)
$\log p$ (pre-duty, excl. charges) 3.40 3.00 1.30–5.40
$\log p$ (pre-duty) 3.50 3.10 1.40–5.40
$\log p^{\text{duty}}$ (post-duty) 3.50 3.10 1.40–5.40
Notes: This table reports summary statistics for the main estimation sample at the importer-supplier-product level. Full details on sample construction are provided in Appendix C.2. Variables are measured at the HS10 product level unless noted otherwise. Panel A reports trade relationship characteristics. $s_{ijh}$ is exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ is buyer $j$’s share in exporter $i$’s U.S. exports of product $h$. Relationship length is the number of years between the supplier-buyer pair’s first transaction in product $h$ and year $t$; Relationship length (all products) is defined analogously, using the pair’s first transaction in any product. # Transactions is the cumulative number of transactions between the pair in product $h$, from their first transaction through year $t$. # Products per pair is the number of distinct HS10 products traded by the pair in year $t$. Buyer tenure (Supplier tenure) is the number of years between the buyer’s (supplier’s) first recorded transaction in the LFTTD and year $t$, pooling across partners and products. Panel B reports log unit values (value divided by quantity). $\log p$ (pre-duty, excl. charges) corresponds to the FOB (Free on Board) unit value, excluding shipping and insurance charges; $\log p$ (pre-duty) includes these charges to obtain the CIF (Cost, Insurance, and Freight) unit value; and $\log p^{\text{duty}}$ (post-duty) further incorporates import duties. Statistics are based on confidential LFTTD data and rounded according to U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Trading relationships are also persistent: the typical importer-exporter pair has been transacting in a given HS10 product for about four years, and in any product for nearly five. Activity within the relationship is intensive: at the median, a pair has accumulated 16 transactions in a product to date and trades 2.5 distinct HS10 products in a given year. Buyer and supplier tenure across all partners and products average about seven and six years, respectively.
Panel B of Table 1 reports descriptive statistics for three bilateral (log) price measures, each defined as transaction value divided by quantity: an FOB (Free on Board) pre-duty price that excludes shipping and insurance charges; a CIF (Cost, Insurance, and Freight) pre-duty price that adds those charges; and a post-duty price that further incorporates import duties. All three are highly dispersed across importer-exporter-product-year observations, with interquartile ranges of about four log points.
@@
To assess the sources of this variation, we conduct a variance decomposition following Fontaine2019 (Appendix D.1). Product-year fixed effects account for about 50% of the total variance, while match-specific residuals account for 4%. Focusing on within supplier–product–year variation, 77% of the remaining dispersion is explained by match-specific factors. This pattern holds across price definitions and underscores the importance of relationship-specific forces in pricing.
3.4 Reduced-Form Evidence on Prices and Pass-Through
As a preliminary exploration of mechanisms, we examine the empirical relationships between markups, pass-through, and bilateral market shares. We interpret these patterns through the lens of Propositions 2 and 4 as descriptive evidence consistent with two-sided market power and decreasing returns to scale, before turning to structural estimation in Section 4.
Markups and Bilateral Market Shares
Proposition 2 implies that under two-sided market power ($\phi\in(0,1)$) and decreasing returns to scale ($\theta<1$), bilateral markups rise in the exporter’s supplier share ($s_{ijht}$) and fall in the importer’s buyer share ($x_{ijht}$). To characterize these empirical patterns, we estimate
@@ 3.4 Reduced-Form Evidence on Prices and Pass-Through
where $\alpha_{s}$ and $\alpha_{x}$ are expected to be positive and negative, respectively. Because markups are not observed, we use log prices as the dependent variable, and include supplier–product–time fixed effects to absorb cost variation and isolate the markup component.
@@
Notes: This table reports OLS and IV estimates of equation (3.1), where the dependent variable is the log CIF duty-excludive unit value of product $h$ imported by buyer $j$ from supplier $i$ in year $t$. Columns alternate between OLS and IV specifications. All regressions control for log relationship length (in years) within HS10 products. Columns (1)–(2) include buyer ($\text{FE}_{j}$), supplier ($\text{FE}_{i}$), and product-year ($\text{FE}_{ht}$) fixed effects. Columns (3)–(4) use supplier–product–year ($\text{FE}_{iht}$) and buyer ($\text{FE}_{j}$) fixed effects. Columns (5)–(6) include fully interacted buyer–product–year and supplier–product–year fixed effects ($\text{FE}_{jht}$, $\text{FE}_{iht}$). IV estimates (even-numbered columns) use leave-one-out instruments: $s_{ijht}$ is instrumented with the average share of other suppliers across buyers of $i$ (excluding $j$), and $x_{ijht}$ with the average share of other buyers across suppliers to $j$ (excluding $i$). Because the model includes multiple endogenous regressors, we report both first-stage and conditional F-statistics from SANDERSON2016212. Standard errors are robust. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
A concern is that market shares mechanically depend on prices, potentially inducing spurious correlation. To address this, we also report IV estimates using leave-one-out measures of market structure. We instrument $s_{ijht}$ with the average supplier share (excluding $i$) across other buyers of exporter $i$ (excluding $j$), and instrument $x_{ijht}$ with the average buyer share (excluding $j$) across other suppliers to importer $j$ (excluding $i$).
Table 2 reports the results. Columns (1)-(2) include exporter, importer, and product-year fixed effects. Columns (3)-(4) replace exporter fixed effects with exporter-product-year fixed effects to absorb marginal costs. Columns (5)-(6) further include importer-product-year fixed effects to capture buyer-specific demand shocks. All specifications control for relationship duration (years since the first shipment of product $h$ between $i$ and $j$). Odd columns report OLS; even columns report IV estimates.
Across specifications, we find $\alpha_{s}>0$ and $\alpha_{x}<0$, both statistically and economically significant. While the magnitudes are not directly interpretable in isolation, these patterns provide clear support for the model’s core mechanisms. Through the lens of Proposition 2, they are consistent with markup variation reflecting both oligopoly and oligopsony forces.
Pass-Through and Decreasing Returns to Scale
Proposition 4 establishes that, in our framework, a negative comovement between pass-through and buyer share arises if and only if production exhibits decreasing returns to scale ($\theta<1$).
We test this implication using the reduced-form regression
$\Delta\ln p^{*}_{ijht}=\$$\alpha_{0}+\alpha_{1}\,\Delta\ln(1+\tau_{cht})+\alpha_{s}\,\Delta\ln(1+\tau_{cht})\cdot s_{ijh,t-1}+\alpha_{x}\,\Delta\ln(1+\tau_{cht})\cdot x_{ijh,t-1}$
$+\alpha_{2}\,s_{ijh,t-1}+\alpha_{3}\,x_{ijh,t-1}+\mathbf{X}_{ijht}\boldsymbol{\gamma}+\mathbf{FE}+\epsilon_{ijht},$ (3.2)
The dependent variable is the change in the duty-exclusive price, denoted $p^{*}_{ijht}$. While the model is expressed in terms of statutory tariffs, observed duty-inclusive prices reflect applied duties and may therefore be subject to measurement error. Focusing on duty-exclusive prices isolates price adjustments net of duties and corresponds to the exporter’s received price in the model. [^22] We write the buyer’s effective price as $p_{ijh}=p^{*}_{ijh}(1+\tau)$, where $p^{*}_{ijh}$ denotes the exporter’s received price and $\tau$ the statutory tariff rate. We focus on the 2017-2018 period, when U.S. imports were subject to large and unexpected tariff increases.
@@
We estimate two specifications. The first includes product-time ($\text{FE}_{ht}$) and exporting country-sector ($\text{FE}_{cs}$) fixed effects, following standard practice in the pass-through literature. This is our baseline. The second is more demanding, combining product–time ($\text{FE}{ht}$) with importer–time ($\text{FE}{jt}$) and exporting country–time ($\text{FE}_{ct}$) fixed effects.
@@
$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.030 0.046
@@
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.029 0.033 0.032
@@
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.407 -0.407 -0.416
@@
Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to CIF duty-exclusive prices at the exporter-importer-product-year level, $\Delta\ln p_{ijht}$. Columns (2) and (6) interact tariffs with the log relationship length (in years) within a product $h$. Columns (3) and (5) interact tariffs with the lagged supplier share, $s_{ijht-1}$, defined as supplier $i$’s share in buyer $j$’s imports of product $h$. Columns (4) and (5) interact tariffs with the lagged buyer share, $x_{ijht-1}$, defined as buyer $j$’s share in supplier $i$’s exports of product $h$. All regressions include product-year and exporter country-sector fixed effects ($\text{FE}_{ht}+\text{FE}_{cs}$). Controls include: (i) $\ln\text{longevity}_{ijht}$; (ii) $\Delta\ln q_{i(-j)ht}$, the change in exporter $i$’s total sales of $h$ to U.S. buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the weighted average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights. Standard errors are clustered at the HS8 product and exporter-country level. The sample corresponds to the ”+ Supplier Multi-Buyer” definition in Table C.1. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).